# Online Episodic Memory Visual Query Localization with Egocentric Streaming Object Memory

### v2

@@ Abstract
[…]
@@ 1 Introduction
Experiments on Ego4D [22] demonstrate the effectiveness of ESOM over other online methods. Results also show that OVQ2D is challenging, with the best method achieving a success rate of only $\sim$4%. This modest performance is mainly due the limitations of current object detection and tracking on real-world egocentric videos. When using oracle detection or tracking, ESOM’s success rate rises to 31.91% and 40.55%, respectively, and reaches 81.92% with both. These results also shed light on the limited performance of current object detection and tracking algorithms when employed in real-world applications, suggesting that more efforts are needed to improve their practical applicability. Unlike standard evaluations that assess these components in isolation, our study offers a competitive benchmark that can be used to assess progress and applicability in real-world downstream tasks.
@@ 4.1 Object Memory
[…]
@@
Interaction with $\mathcal{M}_{ego}$ is achieved through two operations: .$\texttt{read}\texttt{()}$), and .$\texttt{write}\texttt{()}$). The .$\texttt{read}\texttt{()}$) operation is used to access information on specific objects stored in $\mathcal{M}_{ego}$. Following the object-oriented programming notation, we design .read() as an overloaded method that can take as input either an object ID $i$ or a frame ID $t$. $\mathcal{M}_{ego}.\texttt{read}\texttt{(}i\texttt{)}$ returns the list of spatio-temporal localizations related to object instance $\mathbf{O}_{i}$, whereas the $\mathcal{M}_{ego}.\texttt{read}\texttt{(}t\texttt{)}$ operation returns the spatial information of all objects $\mathbf{O}_{i}$ that are memorized at time $t$:
@@
The .$\texttt{write}\texttt{()}$) operation allows to store new object instances or update existing ones. As before, this operation can be called with different input parameters. The syntax
@@
to add $t$, $\mathbf{b}_{i,t}$, $\mathbf{F}_{t}$, and $c_{i,t}$ to object instance $\mathbf{O}_{i}$ at time $t$ (i.e., we update the representation of $\mathbf{O}_{i}$). If $\mathbf{F}_{t}$ is already stored in $\mathcal{M}_{ego}$, no duplicate is stored. When .$\texttt{write}\texttt{()}$) updates an object, it checks for a break from the previous timestamp $\hat{t}$ (i.e., $t-\hat{t}>1$). If occurs, frames $\mathbf{F}_{\bar{t}},\bar{t}\leq\hat{t}$ linked to the prior sequence of contiguous boxes are deleted, unless they belong to another memorized object.
[…]
@@ 4.2 Object Memory Population (OMP)
The OD module is responsible for finding objects of interest in frame $\mathbf{F}_{t}$ which could be later queried by the user. OD is implemented as an object detector [49] which detects a set of $B$ bounding boxes $\mathbf{b}_{k,t}$ of objects in the frame, coupled with confidence scores $s_{k,t}$, with $1\leq k\leq B$. After detecting the objects of interest, the OD module calls $\mathcal{M}_{ego}.\texttt{read}\texttt{(}t\texttt{)}$ to retrieve the bounding boxes of all object instances already present in memory at time $t$. These are objects which may already have been tracked by the OT module at time $t$ from object instances at time $t-1$. The detected boxes are first filtered based on $s_{k,t}>\lambda_{od}$. Those that do not have an IoU $>\lambda_{iou}$ with any of the objects retrieved from the memory, are treated as localizations of new object instances. Filtered $\mathbf{b}_{k,t}$ together with the associated frame $\mathbf{F}_{t}$, are then stored in memory by calling $\mathcal{M}_{ego}$.write($[t,\mathbf{b}_{k,t},\mathbf{F}_{t},c_{k,t}]$), while the others are discarded. See the supplementary material for details.
@@
To enhance the efficiency of the retrieval stage, each call to the .$\texttt{write}\texttt{()}$) method by OT and OD is preceded by a function that computes relevance scores $c_{i,t}$. These scores identify object patches $\phi_{i,t}$, extracted from $\mathbf{b}_{i,t}$ in $\mathbf{F}_{t}$, that contain sufficiently distinctive features for accurate instance recognition. We employ a binary classifier to assess whether each patch $\phi_{i,t}$ meets the quality threshold for effective retrieval (see Supp. material for details). If the classifier determines that a patch is suitable, $c_{i,t}=1$, otherwise, $c_{i,t}=0$.
[…]
@@ 6.2 Comparison with Offline VQ2D
Table 2 shows state-of-the-art offline VQ2D methods. It is important to note that offline VQ2D and our online formulation are not directly comparable, OVQ2D is a much harder task in which methods do not know have access to the queried object in advance; hence these results are reported mainly to discuss the differences between the two approaches. Current offline methods require storing the entire video sequence, resulting in significant memory usageWe consider a fixed memory of 12.1 GB to store a 5-minutes clip as required in the EGO4D benchmark. Longer clips would take more space.. ESOM reduces memory needs by up 7$\times$ (1.7 vs 12.1GB) and increases inference speed by 13$\times$ when compared to VQLoC (3.1 vs 41 s). Memory and computational savings are even more significant when considering the oracular configuration, reported for reference, with 24$\times$ less memory and 80$\times$ increased inference speed. Interestingly, ESOM is competitive even in terms of success rate in oracular settings (81.92 vs 55.89 of VQLoC), which highlights the potential of OVQ2D and ESOM when tracking and detection are solved.
@@ 6.3 Detailed Analysis of ESOM and Ablations
We assessed how the computation of the relevance scores $c_{i,t}$ performed by means of the binary classifier before the .write() process affects the accuracy and speed of QRL (Table 4). In the absence of the retrieval relevance computation, all object representations stored in memory are considered during retrieval, increasing time as more similarity values need to be computed. Downstream OVQ2D performance is also improved by filtering, thanks to the exclusion of suboptimal object representations.
@@
Table 5 presents the performance of ESOM under different configurations of the OD and OT modules, reporting module-specific and VQL task performance metrics: AP $\uparrow$ and Rec $\uparrow$ for object detectionWe merge all detections in a single “object” class., HOTA $\uparrow$ and AssA $\uparrow$ for object tracking [32], and %(s)tAP25, Succ., Size and Time for downstream OVQ2D.
@@
OT modules have a more significant impact on performance. Changing the oracular tracker OToracle with OTegostark while maintaining the oracular object discovery ODoracle brings Succ. down from 81.92 to 31.91. We observe an inverse correlation between downstream performance and general tracking performance. For instance, OTmasa achieves the highest HOTA score (50.1 vs. 21.7), yet yields significantly lower Succ. (9.38 vs. 31.91). This apparent discrepancy can be explained by the nature of the HOTA metric, which is defined as the geometric mean of DetA (detection accuracy within individual frames) and AssA (association accuracy across time) [32]. In our evaluations, DetA tends to be consistently lower than AssA — for example, ODoracle–OTegostark shows 13.7 vs. 34.7, and ODyolo–OTegostark yields 2.5 vs. 15.1—ultimately reducing overall HOTA scores. OVQ2D requires higher AssA than DetA, as accurate object association is essential for constructing coherent and temporally consistent response tracks in $\mathcal{M}_{ego}$. In cases where associations are incorrect, $\mathcal{M}_{ego}$ accumulates fragmented or duplicated tracklets, and DetA penalizes such false positives. ODoracle–OTmasa stands out as an exception, relying on ground truth detections and thus isolating the role of association, which explains its strong HOTA despite limited downstream success. Being explicitly optimized for egovision, OTegostark leads to the best downstream performance.
@@ 7 Conclusions
In the paper, we introduced the novel task of Online Visual Queries 2D (OVQ2D), where a video is processed in a streaming fashion and an object query must be localized without access to the original video. To address this problem, we proposed ESOM (Egocentric Stream Object Memory), a new online approach that detects, tracks, and memorizes essential objects to support efficient visual query retrieval and localization. ESOM significantly optimizes both memory and computational efficiency, reducing memory usage by up to 7$\times$. Additionally, it shortens retrieval times, bringing them down from minutes or tens of seconds to just a few seconds. Experiments on the Ego4D dataset show ESOM’s superior performance when compared with other online methods and its potential when object detection and tracking components are improved. Besides establishing a benchmark for OVQ2D, we also contribute a principled evaluation, based on a real downstream task, for object detection and tracking algorithms for egovision applications.
@@ Acknowledgments.
Progetto PRIN 2022 PNRR - “Tracking in Egovision for Applied Memory (TEAM)” Codice P20225MSER_001. Codici CUP G53D23006680001 (Università di Udine) e E53D23016240001 (Università di Catania). This research has been funded by the European Union, NextGenerationEU – PNRR M4 C2 I1.1 RS Micheloni.
@@ A.1 Object Memory
[…]

### v3

@@ Abstract
Figure 1: Online visual query localization via object memorization and retrieval. We tackle the problem of online episodic memory and propose ESOM (Refer to caption), an architecture that processes an egocentric video ($\mathcal{V}_{ego}$) online and only once to detect ( ), track ( ), and memorize ($\mathcal{M}_{ego}$) user-relevant objects and frames. When retrieving a visual query ($\mathbf{Q}$), ESOM (Refer to caption) searches ( ) its memory ($\mathcal{M}_{ego}$) for the most recent instance where the query was spatio-temporally localized. By avoiding video storage, ESOM optimizes both memory usage and retrieval speed.
@@ 1 Introduction
Experiments on Ego4D [22] demonstrate the effectiveness of ESOM over other online methods. Results also show that OVQ2D is challenging, with the best method achieving a success rate of only $\sim$4%. This modest performance is mainly due to the limitations of current object detection and tracking on real-world egocentric videos. When using oracle detection or tracking, ESOM’s success rate rises to 31.91% and 40.55%, respectively, and reaches 81.92% with both. These results also shed light on the limited performance of current object detection and tracking algorithms when employed in real-world applications, suggesting that more efforts are needed to improve their practical applicability. Unlike standard evaluations that assess these components in isolation, our study offers a competitive benchmark that can be used to assess progress and applicability in real-world downstream tasks.
@@ 4.1 Object Memory
Figure 2: From an egocentric video to a memory of objects. ESOM injects the visual information in each frame $\mathbf{F}_{t}$ of a video $\mathcal{V}_{ego}$ into $\mathcal{M}_{ego}$, an object memory represented as a dynamic list $\mathbf{O}_{i}$ of tuples $\mathbf{o}_{i,t}$ composed of instance-based, frame-level bounding boxes $\mathbf{b}_{i,t}$, related frames $\mathbf{F}_{t}$, and relevance labels $c_{i,t}$. $\mathcal{M}_{ego}$ is built by an Object Memory Population (OMP) algorithm which processes ( Refer to caption) frames $\mathbf{F}_{t}$ online.
@@
Interaction with $\mathcal{M}_{ego}$ is achieved through two operations: $\texttt{read}\texttt{()}$, and $\texttt{write}\texttt{()}.$ The $\texttt{read}\texttt{()}$ operation is used to access information on specific objects stored in $\mathcal{M}_{ego}$. Following the object-oriented programming notation, we design .read() as an overloaded method that can take as input either an object ID $i$ or a frame ID $t$. $\mathcal{M}_{ego}.\texttt{read}\texttt{(}i\texttt{)}$ returns the list of spatio-temporal localizations related to object instance $\mathbf{O}_{i}$, whereas the $\mathcal{M}_{ego}.\texttt{read}\texttt{(}t\texttt{)}$ operation returns the spatial information of all objects $\mathbf{O}_{i}$ that are memorized at time $t$:
@@
The $\texttt{write}\texttt{()}$ operation allows to store new object instances or update existing ones. As before, this operation can be called with different input parameters. The syntax
@@
to add $t$, $\mathbf{b}_{i,t}$, $\mathbf{F}_{t}$, and $c_{i,t}$ to object instance $\mathbf{O}_{i}$ at time $t$ (i.e., we update the representation of $\mathbf{O}_{i}$). If $\mathbf{F}_{t}$ is already stored in $\mathcal{M}_{ego}$, no duplicate is stored. When $\texttt{write}\texttt{()}$ updates an object, it checks for a break from the previous timestamp $\hat{t}$ (i.e., $t-\hat{t}>1$). If occurs, frames $\mathbf{F}_{\bar{t}},\bar{t}\leq\hat{t}$ linked to the prior sequence of contiguous boxes are deleted, unless they belong to another memorized object.
Figure 3: Memory population by object tracking and discovery. The Object Tracking (OT) ( ) reads objects $\mathbf{o}_{i,t-1}$ related to the last frame $\mathbf{F}_{t-1}$ from $\mathcal{M}_{ego}$ and updates their position in current frame $\mathbf{F}_{t}$. In parallel, new objects are detected by the Object Discovery (OD) ( ) module in frame $\mathbf{F}_{t}$. Relevance scores $c_{i,t}$ are computed and $\mathcal{M}_{ego}$ is updated.
@@ 4.2 Object Memory Population (OMP)
The OD module is responsible for finding objects of interest in frame $\mathbf{F}_{t}$ which could be later queried by the user. OD is implemented as an object detector [49] which detects a set of $B$ bounding boxes $\mathbf{b}_{k,t}$ of objects in the frame, coupled with confidence scores $s_{k,t}$, with $1\leq k\leq B$. After detecting the objects of interest, the OD module calls $\mathcal{M}_{ego}.\texttt{read}\texttt{(}t\texttt{)}$ to retrieve the bounding boxes of all object instances already present in memory at time $t$. These are objects which may already have been tracked by the OT module at time $t$ from object instances at time t1t-1. The detected boxes are first filtered based on $s_{k,t}>\lambda_{od}$. Those that do not have an IoU $\lambda_{iou}$ with any of the objects retrieved from the memory, are treated as localizations of new object instances. Filtered $\mathbf{b}_{k,t}$ together with the associated frame $\mathbf{F}_{t}$, are then stored in memory by calling $\mathcal{M}_{ego}$.write($t,\mathbf{b}_{k,t},\mathbf{F}_{t},c_{k,t}]$), while the others are discarded. See the supplementary material for details.
@@
To enhance the efficiency of the retrieval stage, each call to the $\texttt{write}\texttt{()}$ method by OT and OD is preceded by a function that computes relevance scores $c_{i,t}$. These scores identify object patches $\phi_{i,t}$, extracted from $\mathbf{b}_{i,t}$ in $\mathbf{F}_{t}$, that contain sufficiently distinctive features for accurate instance recognition. We employ a binary classifier to assess whether each patch $\phi_{i,t}$ meets the quality threshold for effective retrieval (see Supp. material for details). If the classifier determines that a patch is suitable, $c_{i,t}=1$, otherwise, $c_{i,t}=0$.
Figure 4: Visual query localization by memory retrieval. When the user provides a visual query $\mathbf{Q}$, the Query Retrieval and Localization algorithm ( Refer to caption) is triggered. This algorithm compares the representation of $\mathbf{Q}$ with the representations of each object $\mathbf{O}_{i}$ in $\mathcal{M}_{ego}$. The sequence of contiguous bounding-boxes corresponding to the best-matched ( ) object and the associated RGB frames are retained as the visual response track $\mathbf{r}$ for $\mathbf{Q}$, while all other objects ( ) are discarded.
@@ 6.2 Comparison with Offline VQ2D
[…]
@@ 6.3 Detailed Analysis of ESOM and Ablations
We evaluated the impact of computing relevance scores $c_{i,t}$ via the binary classifier before the .write() step on QRL accuracy and speed (Table 4). Without relevance filtering, all stored object representations are considered during retrieval, increasing computation time as more similarities must be evaluated. Filtering also improves the downstream OVQ2D performance by discarding suboptimal object representations.
@@
Table 5 reports ESOM’s performance across different OD and OT configurations, including module-specific and VQL task metrics: AP $\uparrow$ and Rec $\uparrow$ for object detectionAll detections are merged into a single “object” class., HOTA $\uparrow$ and AssA $\uparrow$ for object tracking [32], and %(s)tAP25, Succ., Size, and Time for downstream OVQ2D.
@@
OT modules have a more significant impact on performance. replacing the oracular tracker OToracle with OTegostark while keeping ODoracle brings Succ. down from 81.92 to 31.91. We observe an inverse correlation between downstream and general tracking performance. For instance, OTmasa achieves the highest HOTA score (50.1 vs. 21.7), yet yields significantly lower Succ. (9.38 vs. 31.91). This stems from HOTA being the geometric mean of DetA (detection accuracy within individual frames) and AssA (association accuracy across time) [32]. In our evaluations, DetA tends to be consistently lower than AssA — for example, ODoracle–OTegostark shows 13.7 vs. 34.7, and ODyolo–OTegostark yields 2.5 vs. 15.1—ultimately reducing overall HOTA scores. OVQ2D requires higher AssA than DetA, as accurate object association is essential for constructing coherent and temporally consistent response tracks in $\mathcal{M}_{ego}$. Incorrect associations lead to fragmented or duplicated tracklets, which DetA penalizes. ODoracle–OTmasa stands out as an exception, relying on ground truth detections and thus isolating the role of association, which explains its strong HOTA despite limited downstream success. Being explicitly optimized for egovision, OTegostark leads to the best downstream performance.
@@ 7 Conclusions
[…]
@@ Acknowledgments.
Progetto PRIN 2022 PNRR - “Tracking in Egovision for Applied Memory (TEAM)” Codice P20225MSER_001. Codici CUP G53D23006680001 (Università di Udine) e E53D23016240001 (Università di Catania). This research has been funded by the European Union, NextGenerationEU – PNRR M4 C2 I1.1 RS Micheloni. MD received funding from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement n. 101151834 PRINNEVOT (CUP G23C24000910006).
@@ A.1 Object Memory
Tracker Detector HOTA $\uparrow$ OWTA $\uparrow$ AssA $\uparrow$ DetA $\uparrow$}